# InterCLIP-MEP: Interactive CLIP and Memory-Enhanced Predictor for Multi-modal Sarcasm Detection

### v4

@@ Abstract
The prevalence of sarcasm in social media, conveyed through text-image combinations, presents significant challenges for sentiment analysis and intention mining. Existing multi-modal sarcasm detection methods have been proven to overestimate performance, as they struggle to effectively capture the intricate sarcastic cues that arise from the interaction between an image and text. To address these issues, we propose InterCLIP-MEP, a novel framework for multi-modal sarcasm detection. Specifically, we introduce an Interactive CLIP (InterCLIP) as the backbone to extract text-image representations, enhancing them by embedding cross-modality information directly within each encoder, thereby improving the representations to capture text-image interactions better. Furthermore, an efficient training strategy is designed to adapt InterCLIP for our proposed Memory-Enhanced Predictor (MEP). MEP uses a dynamic, fixed-length dual-channel memory to store historical knowledge of valuable test samples during inference. It then leverages this memory as a non-parametric classifier to derive the final prediction, offering a more robust recognition of multi-modal sarcasm. Experiments demonstrate that InterCLIP-MEP achieves state-of-the-art performance on the MMSD2.0 benchmark, with an accuracy improvement of 1.08% and an F1 score improvement of 1.51% over the previous best method.
Introduction
Sarcasm, with its inherent subtlety and complexity, plays a significant role in human communication by often conveying irony, mockery, or hidden intentions (Muecke 1982; Gibbs and O’Brien 1991; Gibbs and Colston 2007). The automatic detection of sarcasm from text has been a significant research focus, aiding tasks such as sentiment analysis and intent mining (Pang, Lee et al. 2008; Tsur, Davidov, and Rappoport 2010; Bouazizi and Ohtsuki 2015). With the rapid development of social media platforms like Twitter and Reddit, users frequently use text-image combinations to convey their messages. Therefore, the need for multi-modal sarcasm detection has become increasingly apparent, which presents significant challenges, requiring the system to capture and comprehend the intricate interplay between textual and visual information to identify sarcasm cues.
@@ Introduction
Figure 1: An overview of the shortcomings of existing multi-modal sarcasm detection pipelines. In panels (a) and (b), we present two multi-modal sarcasm detection pipelines, with shortcomings indicated by a red question mark. In panel (c), we visually show an example of multi-modal sarcasm cues correctly or incorrectly recognized in a sarcasm sample.
Despite the notable advancements made by existing studies (Xu, Zeng, and Mao 2020; Pan et al. 2020; Liang et al. 2021, 2022) in this field, Qin et al. (2023) have highlighted that the MMSD benchmark (Cai, Cai, and Wan 2019) employed in these works contains spurious cues. These cues lead to the development of biased models, thereby inflating the perceived effectiveness of these models in detecting sarcasm cues from multi-modal data. By introducing MMSD2.0, a refined benchmark that eliminates spurious cues and corrects mislabeled samples, Qin et al. (2023) demonstrates a substantial performance decline of existing methods when evaluated on MMSD2.0. Although they proposed a promising method that employs a Multi-view CLIP for reliable multi-modal sarcasm detection, it still leaves room for improvement, similar to their previous approaches.
As shown in Figure 1(a) and Figure 1(b), some existing methods (Xu, Zeng, and Mao 2020; Pan et al. 2020; Liang et al. 2021, 2022; Wen, Jia, and Yang 2023; Tian et al. 2023; Wei et al. 2024) use dual unimodal encoders, such as BERT (Devlin et al. 2019) for text and ViT (Dosovitskiy et al. 2021) for image, as the backbone to encode samples. This approach makes it difficult for the model to capture the subtle interactive information between text and image. Although Qin et al. (2023) utilized a multi-modal pre-trained backbone, CLIP (Radford et al. 2021), to encode samples and demonstrated notable results on the MMSD2.0 benchmark, the original CLIP only aligns text and images directly, lacking the nuanced semantic space required for sarcasm detection. To illustrate the identification of multi-modal sarcasm cues, we present in Figure 1(c) an example of a sarcasm sample being correctly and incorrectly recognized. Furthermore, as depicted in Figure 1(a), some approaches (Xu, Zeng, and Mao 2020; Pan et al. 2020; Liang et al. 2021, 2022; Wen, Jia, and Yang 2023; Tian et al. 2023) directly classify sarcastic samples based on their representations. This often leads to high prediction entropy and erroneous outcomes, especially for hard cases with greater uncertainty. To the best of our knowledge, only Wei et al. (2024) considers leveraging historical knowledge of samples to aid in detection. As shown in Figure 1(b), they constructed a static semantic space that stores features of all training samples. When identifying sarcastic samples, they query $k$-nearest neighbors in this semantic space and use a voting mechanism to derive the final predictions. However, relying heavily on a static semantic space may limit its ability to adapt to various test samples.
In this paper, building on these insights, we propose InterCLIP-MEP, a novel framework designed to enhance the reliability of multi-modal sarcasm detection. Our InterCLIP-MEP introduces a refined variant of CLIP, Interactive CLIP (InterCLIP), as the backbone to enhance sample encoding for detecting multi-modal sarcasm. InterCLIP enhances the capture of interactive information between text and image by embedding representations from one modality into the encoder of the other, leading to a more nuanced understanding of multi-modal sarcasm cues. During inference, Our InterCLIP-MEP introduces a Memory-Enhanced Predictor (MEP) to build dynamically a non-parametric classifier from valuable historical features to identify multi-modal sarcasm. Specifically, we design an efficient training strategy to adapt InterCLIP for MEP. Based on the sample representations encoded by InterCLIP, we train a classification module to assign pseudo-labels to samples. Simultaneously, we train a projection module to map samples into a latent space. In this latent space, same-class samples are pulled closer together, while different-class samples are pushed further apart. Additionally, we fine-tune the weights of the self-attention modules in InterCLIP’s text and vision encoders using the low-rank adaptation (LoRA) technique (Hu et al. 2022), building upon the original CLIP weights. This approach allows for more efficient adaptation without significantly increasing the computational burden. Building on InterCLIP and the training strategy, MEP leverages historical knowledge of valuable test samples to identify multi-modal sarcasm. It maintains a dynamic, fixed-length dual-channel memory, with one channel storing historical information of non-sarcastic samples and the other storing information of sarcastic samples. MEP then utilizes this memory as a non-parametric classifier to compute the final prediction for each sample, offering more robust and reliable multi-modal sarcasm detection.
Overall, our contributions can be summarized as follows:
@@
  • We propose InterCLIP-MEP, a novel framework for reliable multi-modal sarcasm detection, which introduces Interactive CLIP (InterCLIP) improving the encoding of text-image interactions by embedding cross-modality information within each encoder and Memory-Enhanced Predictor (MEP) offering more robust and reliable multi-modal sarcasm detection.
  • We develop an efficient training strategy that ensures computational efficiency and superior performance, bridging InterCLIP and MEP.
  • Our extensive experiments on the more reliable benchmark, MMSD2.0, demonstrate that InterCLIP-MEP [^1] https://github.com/CoderChen01/InterCLIP-MEP outperforms existing state-of-the-art methods.
Figure 2: Overview of our framework. Training Interactive CLIP (InterCLIP). Vision and text representations are extracted using separate encoders and embedded into the top-n layers of the opposite modality’s encoder for interaction. The top-n layers are fine-tuned with LoRA, while the rest of the encoder remains frozen. Final vision and text representations are concatenated and used to train a classification module for sarcasm detection. A projection module is also trained to project representations into a latent space. Memory-Enhanced Predictor (MEP). During inference, InterCLIP generates interactive representations. The classification module assigns pseudo-labels, and the projection module provides projection features. MEP updates dynamic memory with these features and pseudo-labels. The final prediction is made by comparing the current sample’s projected feature with those in memory.
Related Work
@@
Multi-modal sarcasm detection. Early work in sarcasm detection primarily focused on textual data (Bouazizi and Ohtsuki 2015; Amir et al. 2016; Baziotis et al. 2018). With the rise of social media, detecting sarcasm from text-image data has become increasingly challenging, leading to the development of several multi-modal approaches. Schifanella et al. (2016) were among the first to use multi-modal posts from social platforms to identify sarcasm cues from both text and images. Cai, Cai, and Wan (2019) introduced the MMSD benchmark for multi-modal sarcasm detection, demonstrating its effectiveness with a hierarchical fusion model incorporating image attributes. This benchmark has been widely used in subsequent studies (Xu, Zeng, and Mao 2020; Pan et al. 2020; Liang et al. 2021, 2022; Liu, Wang, and Li 2022; Qin et al. 2023; Wen, Jia, and Yang 2023; Tian et al. 2023; Wei et al. 2024). However, MMSD has been proven to contain spurious cues that could bias models (Qin et al. 2023). To address this, Qin et al. (2023) introduced the MMSD2.0 benchmark, which removed these cues and corrected mislabeled samples. They observed a significant performance drop when re-evaluating state-of-the-art methods on MMSD2.0 and proposed the need for more stable multi-modal sarcasm detection methods. In this work, we introduce the InterCLIP-MEP framework towards more reliable multi-modal sarcasm detection.
CLIP adaptation. The Contrastive Language-Image Pretraining (CLIP) model (Radford et al. 2021) excels in vision-language tasks. Adapting CLIP for specific domains has shown substantial improvements, as demonstrated by Li, Shakhnarovich, and Yeh (2022) for phrase localization, Liang et al. (2023) for open-vocabulary semantic segmentation, and Wang et al. (2023) for action recognition. In this work, inspired by Ganz et al. (2024), we conditionally enhance both the text and visual encoders of CLIP, making it more effective in capturing the interplay between text and images to identify sarcasm. Unlike Ganz et al. (2024), who focused solely on embedding text into the visual encoder, we also explore embedding images into the text encoder. Furthermore, their approach is limited to general classification tasks and does not address the complexities of multi-modal sarcasm detection.
Memory-enhanced prediction. Inspired by cognitive science (Stokes 2015; Baddeley 2000), memory has been introduced to enhance neural networks (Weston, Chopra, and Bordes 2014; Sukhbaatar et al. 2015). Several studies (Wu et al. 2018; Wen, Jia, and Yang 2023) have used memory mechanisms to improve model training, and some (Zhang et al. 2024; Wei et al. 2024) leverage memory to store historical knowledge, enhancing prediction accuracy. In this work, we introduce a memory-enhanced predictor for multi-modal sarcasm detection. In contrast to other methods, our memory dynamically updates during testing, utilizing relevant historical information for improved accuracy and robustness.
Methodology
@@
Interactive CLIP
The input to Interactive CLIP (InterCLIP) is a text-image pair $\mathcal{P}=(T,I)$, where $T$ represents a piece of text and $I$ represents an image. Here, for simplicity, we do not consider the case of batch inputs. The text encoder $\mathcal{T}$ extracts the vanilla text representation $\mathbf{F}_{t}$:
$\mathbf{F}_{t}=\mathcal{T}(T)=\{h^{t}_{\textrm{bos}}(t_{\textrm{bos}}),h^{t}_{1}(t_{1}),\dots,h^{t}_{n}(t_{n}),h^{t}_{\textrm{eos}}(t_{\textrm{eos}})\},$ (1)
where $t_{i}$ denotes a text token, $n$ is the length of $T$ after tokenization, $t_{\textrm{bos}}$ and $t_{\textrm{eos}}$ are special tokens required by the text encoder. Here, $h^{t}_{i}(\cdot)\in\mathbb{R}^{d_{t}}$ represents the $d_{t}$-dimensional encoded representation of the corresponding token $t_{i}$, with $i$ ranging from 1 to $n$, including the beginning-of-sequence (bos) and end-of-sequence (eos) tokens. The vision encoder $\mathcal{V}$ extracts the vanilla image representation $\mathbf{F}_{v}$:
@@
where $I$ is processed into multiple patches $p_{i}$, $m$ is the number of patches, and $p_{\textrm{cls}}$ is a special token required by the visual encoder. Here, $h^{v}_{i}(\cdot)\in\mathbb{R}^{d_{v}}$ represents the $d_{v}$-dimensional encoded representation of the corresponding $p_{i}$, with $i$ ranging from 1 to $m$, including the classification (cls) token. Specifically, both $\mathbf{F}_{t}$ and $\mathbf{F}_{v}$ are representations from the final layer outputs of their respective encoders. Conditioning on $\mathbf{F}_{t}$ or $\mathbf{F}_{v}$, we can obtain the interactive text representation $\tilde{\mathbf{F}}_{t}$ or the interactive image representation $\tilde{\mathbf{F}}_{v}$:
@@
To be specific, we condition only the top-$n$ self-attention layers of the text or vision encoder, where $n$ is a hyperparameter that will be analyzed in the experiment section. Figure 3 illustrates the structure of the conditioned self-attention layers. Given that the text and vision encoder in CLIP share a similar architecture, for simplicity, we denote the input representations to the self-attention layers of the text or vision encoder as $\mathbf{H}_{t/v}$, which are derived from the outputs of the previous layer. The previous layer can either be conditioned or non-conditioned. Due to the dimensional mismatch between the embedded representations $\mathbf{F}_{v/t}$ and the corresponding encoder representation space, we introduce an adapting projection layer $\mathcal{F}_{t/v}$ to project $\mathbf{F}_{v/t}$ into the appropriate representation space.
@@
To fuse the input representations $\mathbf{H}_{t/v}$ with the projected embedded representations $\mathbf{F}_{v/t}^{{}^{\prime}}=\mathcal{F}_{t/v}(\mathbf{F}_{v/t})$, we concatenate them and feed them into the attention layer to obtain the transformed representations. We then extract the transformed input representations $\mathbf{H}^{{}^{\prime}}_{t/v}$ from the output. Following Ganz et al. (2024), we apply a gated projection layer $\mathcal{G}_{t/v}$ along with the self-attention’s projection head $\mathcal{H}_{t/v}$ using a learnable gating mechanism to compute the self-attention output representation $\mathbf{H}^{{}^{\prime\prime}}_{t/v}$.
Given the similarity between the self-attention layers of the vision encoder and the text encoder, we use the text encoder $\mathcal{T}$ to illustrate the process as follows:
@@
$\begin{split}\mathbf{F}_{v}^{{}^{\prime}}&=\mathcal{F}_{t}(\mathbf{F}_{v}),\quad\mathbf{F}_{v}\in\mathbb{R}^{m\times d_{v}},\quad\mathbf{F}_{v}^{{}^{\prime}}\in\mathbb{R}^{m\times d_{t}},\\ \mathbf{H}^{{}^{\prime}}_{t}&=\textrm{Attn}_{t}(\mathbf{H}_{t}\oplus\mathbf{F}_{v}^{{}^{\prime}})_{[:n]},\\ &\phantom{=}\mathbf{H}_{t},\mathbf{H}^{{}^{\prime}}_{t}\in\mathbb{R}^{n\times d_{t}},\quad\mathbf{H}_{t}\oplus\mathbf{F}_{v}^{{}^{\prime}}\in\mathbb{R}^{(n+m)\times d_{t}},\\ \mathbf{H}^{{}^{\prime\prime}}_{t}&=\mathcal{H}_{t}(\mathbf{H}^{{}^{\prime}}_{t})+\mathcal{G}_{t}(\mathbf{H}^{{}^{\prime}}_{t})\cdot\textrm{tanh}(\beta_{t}),\quad\mathbf{H}^{{}^{\prime\prime}}_{t}\in\mathbb{R}^{n\times d_{t}}.\end{split}$ (4)
@@
Here, $\oplus$ denotes the concatenation operation, and $\beta_{t}$ is a learnable parameter initialized to 0 to ensure training stability. The subsequent computation follows the original CLIP (Radford et al. 2021), ultimately yielding the interactive representations $\tilde{\mathbf{F}}_{t}$.
Figure 3: Structure of the conditional self-attention layer.
InterCLIP has three distinct modes to achieve the final fused feature $\tilde{h}^{f}\in\mathbb{R}^{d_{t}+d_{v}}$:
  • T2V: The text encoder extracts the vanilla text representation $\mathbf{F}_{t}$. This text representation is then embedded into the vision encoder to obtain the conditioned vision representation $\tilde{\mathbf{F}}_{v}$. The final fused representation $\tilde{h}^{f}$ is formed by concatenating $h^{t}_{\textrm{eos}}(t_{\textrm{eos}})$ and $\tilde{h}^{v}_{\textrm{cls}}(p_{\textrm{cls}})$.
@@
  • V2T: The vision encoder extracts the vanilla image representation $\mathbf{F}_{v}$. This image representation is then embedded into the text encoder to obtain the conditioned text representation $\tilde{\mathbf{F}}_{t}$. The final fused representation $\tilde{h}^{f}$ is formed by concatenating $\tilde{h}^{t}_{\textrm{eos}}(t_{\textrm{eos}})$ and $h^{v}_{\textrm{cls}}(p_{\textrm{cls}})$.
@@
  • Two-way (TW): The text and vision encoders extract the vanilla text and image representations $\mathbf{F}_{t}$ and $\mathbf{F}_{v}$, respectively. These representations are then embedded into the vision and text encoders to obtain the conditioned image representation $\tilde{\mathbf{F}}_{v}$ and text representation $\tilde{\mathbf{F}}_{t}$. The final fused representation $\tilde{h}^{f}$ is formed by concatenating $\tilde{h}^{t}_{\textrm{eos}}(t_{\textrm{eos}})$ and $\tilde{h}^{v}_{\textrm{cls}}(p_{\textrm{cls}})$.
@@
We will analyze the effectiveness of these three modes of interaction in the experimental analysis.
Training Strategy
@@
The classification module $\mathcal{F}_{c}$ calculates the probability of a sample being sarcastic or non-sarcastic:
$\hat{y}=\textrm{softmax}(\mathcal{F}_{c}(\tilde{H}^{f})),$ (5)
where $\tilde{H}^{f}\in\mathbb{R}^{{N}\times(d_{t}+d_{v})}$ represents the fused representations of a batch of samples, and $N$ denotes the batch size. We optimize $\mathcal{F}_{c}$ using binary cross-entropy loss:
$\mathcal{L}^{c}=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log(\hat{y}_{i,1})+(1-y_{i})\log(1-\hat{y}_{i,0})\right],$ (6)
@@
where $\hat{H}^{f}_{P}$ and $\hat{H}^{f}_{N}$ represent the projected fused features of positive and negative samples, respectively.
We fully train the modules $\mathcal{F}_{c}$, $\mathcal{F}_{p}$, the adapting projection layer $\mathcal{F}_{t/v}$, the gated projection layer $\mathcal{G}_{t/v}$, and the parameters $\beta_{t/v}$. We use LoRA (Hu et al. 2022) to fine-tune parts of the weight matrices $\mathbf{W}$ in the self-attention modules of all encoders, specifically various combinations of $W_{q}$, $W_{k}$, $W_{v}$, and $W_{o}$. We consider $\mathbf{W}$ and the rank $r$ of LoRA as hyperparameters for our study. All modules are optimized by minimizing the joint loss:
@@
Memory-Enhanced Predictor
As depicted in Figure 2 (right), we present the Memory-Enhanced Predictor (MEP) that builds upon the learned InterCLIP, along with the classification module and the projection module, leveraging the valuable historical knowledge of test samples to enhance sarcasm detection.

### v5

@@ Abstract
Sarcasm in social media, often expressed through text-image combinations, poses challenges for sentiment analysis and intention mining. Current multi-modal sarcasm detection methods have been demonstrated to overly rely on spurious cues within the textual modality, revealing a limited ability to genuinely identify sarcasm through nuanced text-image interactions. To solve this problem, we propose InterCLIP-MEP, which introduces Interactive CLIP (InterCLIP) with an efficient training strategy to extract enriched text-image representations by embedding cross-modal information directly into each encoder. Additionally, we design a Memory-Enhanced Predictor (MEP) with a dynamic dual-channel memory that stores valuable test sample knowledge during inference, acting as a non-parametric classifier for robust sarcasm recognition. Experiments on two benchmarks demonstrate that InterCLIP-MEP achieves state-of-the-art performance, with significant accuracy and F1 score improvements on MMSD and MMSD2.0. Our code is available at https://github.com/CoderChen01/InterCLIP-MEP.
InterCLIP-MEP: Interactive CLIP and Memory-Enhanced Predictor for Multi-modal Sarcasm Detection
Junjie Chen1,3thanks: Junjie is a research assistant at the EPIC Lab of Shanghai Jiao Tong University, working remotely with Linfeng Zhang.    Hang Yu2thanks: Co-corresponding author.    Subin Huang3thanks: Corresponding author.    Sanmin Liu3    Linfeng Zhang1 1Shanghai Jiao Tong University, China    2Shanghai University, China 3Anhui Polytechnic University, China jorji.chen@gmail.com,  yuhang@shu.edu.cn,  subinhuang@ahpu.edu.cn, sanmin.liu@ahpu.edu.cn,  zhanglinfeng@sjtu.edu.cn
1 Introduction
Sarcasm, with its subtlety and complexity, plays a key role in communication by conveying irony, mockery, or hidden meanings Muecke (1982); Gibbs and O’Brien (1991); Gibbs and Colston (2007). Automatically detecting sarcasm in text has become an important research area, supporting tasks like sentiment analysis and intent mining Pang et al. (2008); Tsur et al. (2010); Bouazizi and Ohtsuki (2015).
@@ 1 Introduction
Figure 1: An overview of the shortcomings of existing multi-modal sarcasm detection pipelines. In panels (a) and (b), we present two main multi-modal sarcasm detection pipelines, with shortcomings indicated by a red question mark. In panel (c), we visually show an example of multi-modal sarcasm cues correctly or incorrectly recognized in a multi-modal sarcasm sample.
With the rise of social media platforms like Twitter and Reddit, users often use text-image combinations to express their messages. As a result, multi-modal sarcasm detection is increasingly important, posing challenges in understanding the complex relationship between textual and visual cues to identify sarcasm.
As shown in Figures 1(a) and 1(b), many methods rely on dual unimodal pre-trained encoders, such as ViTDosovitskiy et al. (2021) and BERT Devlin et al. (2019), as the backbone for encoding text-image pairs, followed by specific feature fusion Xu et al. (2020); Pan et al. (2020); Liang et al. (2021, 2022); Wen et al. (2023); Tian et al. (2023); Wei et al. (2024). However, this approach may not capture multi-modal sarcasm cues as effectively as multi-modal pre-trained models like CLIP Radford et al. (2021). In Figure 1(a), the use of a learnable classification head to predict labels from fused representations is common but often associated with high predictive entropy and significant uncertainty. Wei et al. (2024) pioneered the construction of a static semantic space using historical training samples, where more robust predictions are obtained during inference through KNN-based querying and voting, as illustrated in Figure 1(b). However, while CLIP has been proven to be effective in serving as a text-image encoder for multi-modal sarcasm detection Qin et al. (2023), it still struggles to capture multi-modal sarcasm cues due to the inherent inconsistency of sarcasm, which conflicts with CLIP’s direct alignment of text and image. Furthermore, relying on a static semantic space for inference is ill-suited to handle the dynamic nature of evolving sample distributions. In fact, Qin et al. (2023) have shown that many models rely on spurious cues in the MMSD benchmark Cai et al. (2019), leading to biased results.
Building on the limitations of prior multi-modal sarcasm detection approaches, we propose Interactive CLIP (InterCLIP) as the backbone, embedding cross-modal representations directly into text and vision encoders to enhance the understanding of multi-modal sarcasm cues (Figure 2, left). To complement this, we design a Memory-Enhanced Predictor (MEP) that dynamically utilizes historical test sample features to create a more adaptive and reliable non-parametric classifier for final predictions (Figure 2, right). Together, these components form the proposed framework, InterCLIP-MEP. Furthermore, InterCLIP-MEP employs an efficient training strategy that fine-tunes cross-modal interactions through a lightweight adaptation mechanism, ensuring computational efficiency while delivering state-of-the-art performance (Figure 2, left). Overall, our contributions are as follows:
@@
  • We introduce InterCLIP-MEP [^1] Our code is available in the supplementary material., a novel framework for multi-modal sarcasm detection, which combines Interactive CLIP (InterCLIP) for enhanced text-image interaction encoding and Memory-Enhanced Predictor (MEP) for more robust and reliable sarcasm predictions.
  • We propose an efficient training strategy that significantly reduces computational overhead compared to state-of-the-art methods. By introducing approximately 20.6x fewer trainable parameters, our approach reduces GPU memory usage by about 2.5x and accelerates computation by roughly 8.7x with a batch size of 128, all while maintaining superior performance on a single NVIDIA RTX 4090 GPU.
  • Through extensive experiments on the MMSD and MMSD2.0 benchmarks, we show that InterCLIP-MEP improves accuracy by 1.08% and F1 score by 1.51% over state-of-the-art methods, especially on MMSD2.0.
2 Related Work
@@
Early research in sarcasm detection focused primarily on text data Bouazizi and Ohtsuki (2015); Amir et al. (2016); Baziotis et al. (2018). With the rise of social media, detecting sarcasm in text-image pairs has become more challenging, driving the development of multi-modal approaches. Schifanella et al. (2016) were among the first to explore multi-modal social media posts for identifying sarcasm cues. Building on this, Cai et al. (2019) introduced the MMSD benchmark, demonstrating the effectiveness of a hierarchical fusion model that integrates image features. This benchmark has since become a foundation for multi-modal sarcasm detection, inspiring subsequent studies Xu et al. (2020); Pan et al. (2020); Liang et al. (2021, 2022); Liu et al. (2022); Qin et al. (2023); Wen et al. (2023); Tian et al. (2023); Wei et al. (2024).
However, the MMSD benchmark was later found to contain spurious cues that could lead to model bias Qin et al. (2023). To mitigate this, Qin et al. (2023) introduced the MMSD2.0 benchmark, which removes these cues and corrects mislabeled samples. Re-evaluations on MMSD2.0 revealed significant performance drops in existing models, emphasizing the need for more robust approaches. In parallel, Tang et al. (2024) explored the use of large language models (LLMs) in multi-modal sarcasm detection, incorporating instruction templates and retrieval modules. While promising, the performance improvements were modest compared to the substantial computational cost.
In this work, we present InterCLIP-MEP, a lightweight and efficient framework that achieves competitive performance without the high resource demands of LLM-based approaches. By overcoming the limitations of current methods, our approach offers a practical and scalable solution for multi-modal sarcasm detection.
Figure 2: Overview of our framework. (I) Training Interactive CLIP (InterCLIP): Vision and text representations are extracted using separate encoders and embedded into the top-$n$ layers of the opposite modality’s encoder for interaction. The top-$n$ layers are fine-tuned with LoRA, while the rest of the encoder remains frozen. Final vision and text representations are concatenated and used to train a classification module for identifying multi-modal sarcasm. A projection module is also trained to project representations into a latent space. (II) Memory-Enhanced Predictor (MEP): During inference, InterCLIP generates interactive representations. The classification module assigns pseudo-labels, and the projection module provides projection features. MEP updates dynamic memory with these features and pseudo-labels. The final prediction of the current sample is made by comparing its projected feature with those in memory.
3 Methodology
@@
3.1 Interactive CLIP
The input to Interactive CLIP (InterCLIP) is a text-image pair $\mathcal{P}=(T,I)$, where $T$ represents a piece of text and $I$ represents an image. Here, for simplicity, we do not consider the case of batch inputs.
The text encoder $\mathcal{T}$ extracts the vanilla text representations $\mathbf{F}_{t}$:
$\mathbf{F}_{t}$ $=\mathcal{T}(T)$
=$=\{h^{t}_{\textrm{bos}}(t_{\textrm{bos}}),h^{t}_{1}(t_{1}),\dots,h^{t}_{n}(t_{n}),h^{t}_{\textrm{eos}}(t_{\textrm{eos}})\},$ (1)
where $t_{i}$ denotes a text token, $n$ is the length of $T$ after tokenization, $t_{\textrm{bos}}$ and $t_{\textrm{eos}}$ are special tokens required by the text encoder. Here, $h^{t}_{i}(\cdot)\in\mathbb{R}^{d_{t}}$ represents the $d_{t}$-dimensional encoded representation of the corresponding token $t_{i}$, with $i$ ranging from 1 to $n$, including the beginning-of-sequence (bos) and end-of-sequence (eos) tokens.
The vision encoder $\mathcal{V}$ extracts the vanilla image representations $\mathbf{F}_{v}$:
@@ 3.1 Interactive CLIP
where $I$ is processed into multiple patches $p_{i}$, $m$ is the number of patches, and $p_{\textrm{cls}}$ is a special token required by the visual encoder. Here, $h^{v}_{i}(\cdot)\in\mathbb{R}^{d_{v}}$ represents the $d_{v}$-dimensional encoded representation of the corresponding $p_{i}$, with $i$ ranging from 1 to $m$, including the classification (cls) token. Specifically, both $\mathbf{F}_{t}$ and $\mathbf{F}_{v}$ are representations from the final layer outputs of their respective encoders. Conditioning on $\mathbf{F}_{t}$ or $\mathbf{F}_{v}$, we can obtain the interactive text representations $\tilde{\mathbf{F}}_{t}$ or the interactive image representations $\tilde{\mathbf{F}}_{v}$:
@@
To be specific, we condition only the top-$n$ self-attention layers of the text or vision encoder, where $n$ is a hyperparameter that will be analyzed in the experiment section. Figure 3 illustrates the structure of the conditioned self-attention layers. Given that the text and vision encoder in CLIP share a similar architecture, for brevity, we denote the input representations to the self-attention layers of the text or vision encoder as $\mathbf{H}_{t/v}$, which are derived from the outputs of the previous layer. The previous layer can either be conditioned or non-conditioned. Due to the dimensional mismatch between the embedded representations $\mathbf{F}_{v/t}$ and the corresponding encoder representation space, we introduce an adapting projection layer $\mathcal{F}_{t/v}$ to project $\mathbf{F}_{v/t}$ into the appropriate representation space.
@@
To fuse the input representations $\mathbf{H}_{t/v}$ with the projected embedded representations $\mathbf{F}_{v/t}^{{}^{\prime}}=\mathcal{F}_{t/v}(\mathbf{F}_{v/t})$, we concatenate them and feed them into the attention layer to obtain the transformed representations. We then extract the transformed input representations $\mathbf{H}^{{}^{\prime}}_{t/v}$ from the output. Following Ganz et al. (2024), we apply a gated projection layer $\mathcal{G}_{t/v}$ along with the self-attention’s projection head $\mathcal{H}_{t/v}$ using a learnable gating mechanism to compute the self-attention output representations $\mathbf{H}^{{}^{\prime\prime}}_{t/v}$. Given the similarity between the self-attention layers of the vision encoder and the text encoder, we use the text encoder $\mathcal{T}$ to illustrate the process as follows:
@@
$\begin{split}\mathbf{F}_{v}^{{}^{\prime}}&=\mathcal{F}_{t}(\mathbf{F}_{v}),\quad\mathbf{F}_{v}\in\mathbb{R}^{m\times d_{v}},\mathbf{F}_{v}^{{}^{\prime}}\in\mathbb{R}^{m\times d_{t}},\\ \mathbf{H}^{{}^{\prime}}_{t}&=\textrm{Attn}_{t}(\mathbf{H}_{t}\oplus\mathbf{F}_{v}^{{}^{\prime}})_{[:n]},\\ &\phantom{=}\mathbf{H}_{t},\mathbf{H}^{{}^{\prime}}_{t}\in\mathbb{R}^{n\times d_{t}},\mathbf{H}_{t}\oplus\mathbf{F}_{v}^{{}^{\prime}}\in\mathbb{R}^{(n+m)\times d_{t}},\\ \mathbf{H}^{{}^{\prime\prime}}_{t}&=\mathcal{H}_{t}(\mathbf{H}^{{}^{\prime}}_{t})+\mathcal{G}_{t}(\mathbf{H}^{{}^{\prime}}_{t})\cdot\textrm{tanh}(\beta_{t}),\mathbf{H}^{{}^{\prime\prime}}_{t}\in\mathbb{R}^{n\times d_{t}}.\end{split}$ (4)
@@
Here, $\oplus$ denotes the concatenation operation, and $\beta_{t}$ is a learnable gating parameter initialized to 0 to ensure training stability. The subsequent computation follows the original CLIP Radford et al. (2021), ultimately yielding the interactive representations $\tilde{\mathbf{F}}_{t}$.
Figure 3: Structure of the conditional self-attention.
InterCLIP supports three interaction modes for fusing text and image features into the final representation $\tilde{h}^{f}\in\mathbb{R}^{d_{t}+d_{v}}$:
@@
  • T2V: Text representations $\mathbf{F}_{t}$ are embedded into the vision encoder to produce interactive image representations $\tilde{\mathbf{F}}_{v}$. $\tilde{h}^{f}$ is formed by concatenating $h^{t}_{\textrm{eos}}$ and $\tilde{h}^{v}_{\textrm{cls}}$.
@@
  • V2T: Image representations $\mathbf{F}_{v}$ are embedded into the text encoder to produce interactive text representations $\tilde{\mathbf{F}}_{t}$. $\tilde{h}^{f}$ is formed by concatenating $\tilde{h}^{t}_{\textrm{eos}}$ and $h^{v}_{\textrm{cls}}$.
  • Two-way (TW): Both text and image representations $\mathbf{F}_{t}$ and $\mathbf{F}_{v}$ are embedded into each other’s encoders, resulting in $\tilde{\mathbf{F}}_{t}$ and $\tilde{\mathbf{F}}_{v}$. $\tilde{h}^{f}$ is formed by concatenating $\tilde{h}^{t}_{\textrm{eos}}$ and $\tilde{h}^{v}_{\textrm{cls}}$.
@@
We will analyze the effectiveness of these three interaction modes in the experimental analysis.
Training Strategy.
@@
Given the fused features of a batch of samples $\tilde{H}^{f}\in\mathbb{R}^{{N}\times(d_{t}+d_{v})}$, the classification module $\mathcal{F}_{c}$ calculates the probabilities $\hat{y}$ of these samples being sarcastic or non-sarcastic:
$\hat{y}=\textrm{softmax}(\mathcal{F}_{c}(\tilde{H}^{f})),\quad\hat{y}\in\mathbb{R}^{N\times 2},$ (5)
where $N$ denotes the batch size. We optimize $\mathcal{F}_{c}$ using binary cross-entropy loss:
$\mathcal{L}^{c}=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log(\hat{y}_{i,1})+(1-y_{i})\log(1-\hat{y}_{i,0})\right],$
@@
where $\hat{H}^{f}_{P}$ and $\hat{H}^{f}_{N}$ represent the projected features of positive and negative samples, respectively.
We fully train the modules $\mathcal{F}_{c}$, $\mathcal{F}_{p}$, the adapting projection layers ($\mathcal{F}_{t}$ and $\mathcal{F}_{v}$), the gated projection layers ($\mathcal{G}_{t}$ and $\mathcal{G}_{v}$), and the learnable gating parameters ($\beta_{t}$ and $\beta_{v}$). We use LoRA Hu et al. (2022) to fine-tune parts of the weight matrices $\mathbf{W}$ in the self-attention modules of all encoders, specifically various combinations of $W_{q}$, $W_{k}$, $W_{v}$, and $W_{o}$. We consider $\mathbf{W}$ and the rank $r$ of LoRA as hyperparameters for our study. All learnable parts are optimized by minimizing the joint loss: